[Search for users] [Overall Top Noters] [List of all Conferences] [Download this site]

Conference asdg::cmp

Title:CMP
Notice:Welcome to the CMP notesfile
Moderator:FABSIX::B_TOWERLLE
Created:Tue Oct 31 1995
Last Modified:Fri Jun 06 1997
Last Successful Update:Fri Jun 06 1997
Number of topics:94
Total number of notes:6381

31.0. "SPEEDFAM GENERAL ISSUES" by LUDWIG::TDYER () Wed Jul 03 1996 12:43

    this is for genral SPEEDFAM issues, procedures, contacts, etc.
T.RTitleUserPersonal
Name
DateLines
31.1COND DF PROCEDURELUDWIG::TDYERWed Jul 03 1996 12:4728

 CONDITIONER DOWN FORCE CALIBRATION PROCEDURE: (7/3/96 TPD)
____________________________________________________________________________

 1) GO TO THE MANUAL CONDITIONER SCREEN
	a. Set osc ID and OD to 119, press "conditioner to osc OD"
 2) REMOVE CONDITIONER RING AND BLACK RETAINING RING, PLACE ONTOP OF 
    THE END OF THE CONDITIONER ARM (for correct weight).
 3) PLACE DOWN FORCE CALIBRATION PLATE UNDER THE SILVER 6 BOLT RETAINING
    RING.
 4) GO TO THE DIAGNOSTIC, MANUAL DIGITAL OUTPUT PAGE
	a. Turn off DG-OUT 54.
 5) GO TO THE MANUAL ANALOG PAGE.
	a. Select AN-OUT 5.
	b. Select 4095
 	c. Press set value
 6) ADJUST THE GAIN POT ON ER6 (electronic pressure regulator) IN THE
    PNEUMATIC CABINET UNTIL THE DOWN FORCE METER READS 77lbs.
 7) ON THE ANALOG PAGE SET AO-5 TO ZERO, THAN BACK TO 4095, RE-CALIBRATE
    THE GAIN POT AS NEEDED.
 8) SET AO-5 TO ZERO
 9) GO TO THE MANUAL CONDITIONER PAGE AND PRESS COND ARM UP
10) INSTALL CONDITIONER RETAINING RING AND CONDITIONER WHEEL.
11) PRESS CONDITIONER TO HOME.
12) CALIBRATION IS COMPLETE, PUT SYSTEM IN WET MODE.


31.2SF REPORTING PROC/##SLUDWIG::TDYERWed Jul 03 1996 12:5744
                                                    
    +---+---+---+---+---+---+---+
    | d | i | g | i | t | a | l |	  
    +---+---+---+---+---+---+tm-+
        Semiconductor
					I N T E R O F F I C E   M E M O

     To: CMP.dis		               Date: JULY 3RD 1996
         		                       From: Timothy P. Dyer
                                               Dept: Engineering
     cc: Dennis Goodwin 	               Ext:  5072  
         Kevin Colman	                       Node: Scoman::Tdyer
         Richard Barrett
                            
==============================================================================
 SUBJECT: SPEEDFAM DOWN RESPONSE AND ESCALATION PROCEDURE
++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++
                                      
    MFG and EEG Techs, please follow this procedure for ANY UNSCHEDULED DOWN 
    events on the CMP.B tools.
 
 MONDAY - FRIDAY 8am - 5pm: MFG PAGE ONE DEC EEG TECH, THEY WILL PAGE THE
			    ONSITE SF REP AT 361.
  
 AFTER 5pm AND WEEKENDS: AFTER 4 HOURS OF DOWN TIME OR IF YOU NEED
    			 TECHNICAL SUPPORT, PAGE THE NEW SPEEDFAM 24 HOUR
                         TECHNICAL RESPONCE CENTER AT 800-276-9149. 
                         THE CENTER WILL CONTACT THE ON CALL TECH SUP ENG 
                         AND OUR LOCAL REPS.
                              
     The reps should not be paged directly after 8pm and weekends as per 
     contract. 
           KEVIN 1-800-796-7363 pin 1020584
 	   RICH  1-800-796-7363 pin 1043322
 
 IF YOU GET UNSATISFACTORYRESPONSE FROM THE 24 LINE OR REQUIRE MORE TECHNICAL
 ASSISTANCE I CAN BE REACHED AT HOME (508-699-4905) OR PAGED (508-387-1389).

 IF THE SYSTEMS ARE DOWN FOR AN EXTENDED TIME OVER THE WEEKENDS ~12+ HOURS, 
 PLEASE GIVE ME A PAGE or CALL REGARDLESS IF SF IS SUPPORTING YOU OR NOT.
 This is flexible, use your judgment, if 4:00am is the 12 hour mark, you 
 can wait till sunrise....please.

 Please file and save this note.......thanks, Tim
31.3SPEEDFAM SCRUBBER RECIPESTRATA::KRUPAThu Aug 01 1996 18:3611
    Here is the scrubber recipes for the Speedfam tools.  If there is ever
    a question about the recipe on the tool, refer to this note.
    
    time = 7 sec for B2&B3; 11 sec for B1
    brush rpm = 300
    brush pressure = 30 psi
    CCW
    Water rinse ON
    
    thanks,
    -Frank
31.4slurry lossLUDWIG::RIDLONEliminate the obliviousSun Sep 29 1996 05:004
    Tools spent an hour down due to a loss of slurry to the room. 
    
    
    						B-shift
31.5no slurrySTRATA::RIDLONEliminate the obliviousTue Oct 01 1996 05:4214
    Lost slurry to the room at around 2345 (caused B2's crash). We had
    originally thought that the crash was due to the breakin wafers being
    thin so we didn't notice the slurry problem until around 0030. Luckily
    all tools were down so nothing was being run. AGAIN THE BLUE SLURRY
    PUMP ALARM DIDN'T GO OFF. Ashland came in and got slurry back up at
    0230. He said the filter wasn't clogged and that all he did was switch
    to pump 2 and flow came up. Flowed the slurry on b3 and b1 and it
    looked good. 
    
    				
    						B-shift
    
    
    
31.6particlesSTRATA::RIDLONEliminate the obliviousTue Oct 22 1996 00:189
    All tools failed for particles. Tim Dyer called into the room and
    requested that the day tank be flushed and cleaned, the filter changed,
    and a new slurry lot number used. Ashland was notified and could do the
    first two but we don't have any new lot numbers in house. Once he was
    done we purged slurry through the lines for about 5 minutes and then di
    flushed for 10. The pads were then scrubbed and mqc's requested.
    
    
    					B-shift
31.7filter changeSTRATA::FINNFri Oct 25 1996 17:3729
The Fab6 CMP slurry system has been reconnected to the Osmmonic Filter 
system. Ashland Chemical has agreed to replace the filters twice this weekend 
to preventive filter clogging and the subsequent loss of slurry. Due to the 
current filter configuration, it will be necessary to shut the CMP room 
DOWN for a period of approximately one hour during each of these two filter 
change events. 

	The First filter change will occur tomorrow at 4:00PM. 

	The Second filter change will occur Monday morning at 6:30 AM.

Please ensure that all processing has stopped prior to these times. In 
addition Ashland Technicians will call the room at ext:4774 to confirm that 
processing on the tools has ceased prior to their beginning work.

Ashland Technicians will again call the room when their work has been
completed and we may continue processing. Their primary contact will be the 
operation group at ext 4774. In the event they are unable to reach anyone in 
the room, the following names have been given to Ashland as backup contacts:

	 John Swenson (pager 697) for Saturday afternoon 4:00pm
	 Mitch Ridlon (pager 652) for Monday morning 6:30 am

Thank you for your anticipated cooperation. 


Brendan 

31.8filter good so farLUDWIG::RIDLONEliminate the obliviousSun Oct 27 1996 21:042
    Slurry room checked at 2015. Input pressure to filter was 26 output was
    22. Flow good, pumping good. Will keep an eye on it.
31.9filter checkLUDWIG::RIDLONEliminate the obliviousMon Oct 28 1996 04:112
    Checked slurry room at 0245. Pressure holding steady at 26 in 22 out. 
    Flow looking good and pumping sounds steady. 
31.10filter checkLUDWIG::RIDLONEliminate the obliviousMon Oct 28 1996 05:502
    Checked slurry room at 0430. Pressure maintaining at 26 in 22 out.
    Flow's good, pump still pumping. 
31.11filter updateLUDWIG::RIDLONEliminate the obliviousMon Oct 28 1996 06:384
    Checked slurry room at 0630. Pressure starting to fall off. 24 in 19
    out. Flow's still good. Pump starting to work harder. Talked with
    Ashland dudes; eight filters due in some time today. Looks like we made
    it through the night.
31.12Cmp.B3 ionizer turned off FABSIX::J_SWENSONFri Nov 08 1996 17:566
     The Ionizer in B3 has been turned off. If any mapping errors occur
     please record them. 
    
    
    
                                   C-SHift
31.14filters/rebuild kits for speedfam vac. FABSIX::J_SWENSONSat Jan 11 1997 18:0713
    Vacuum pump rebuild kits
    ------------------------
    
    
       The kits are located in the blue vidmar over by b4, the diaphrams
       are on back order and will be here within a couple of weeks.
    
       Also the elements for the vacuum line filters are in the same draw.
    
       The draw is also labeled.
    
    
                              C-shift 
31.15Installing/breakin of new conditionersSTRATA::WOOLDRIDGEPleasure, Spiked With PainMon Jan 13 1997 13:5932
    This is the best known new conditioning wheel procedure for the 
speedfams:
          
Rev-1 
=====

Subject: Conditioner ring change procedure

Summary: To install a new conditioner ring that has higher cut-rate in an
         effort to increase removal rates.  

Plan:	- Do not remove pad
	- Check pad step height to ensure there is >60 mils if it isn't 
	  install new pad
	- Remove old ring and save it for Kyle labeled "CMP.x's old ring"
	- Install New conditioner ring w/ cut of 16 to 18 from stores.
	- Run conditioner manually at 80% downforce for 30 minutes on old pad
	- Change pad
	- Record total number of wafers on the tool(Note this is the total
	  running counter that is not re-settable). This is being done to 
	  track the total # of wafers on the conditioner ring.  Record this
	  number and serial# in notes with a comment that a new ring was 
	  installed.
	- Perform std 10 minute condition followed by normal breakins and
	  MQCs.
	- If removal rate has increased be sure to pass it down to next shift.
        - Condioning rings from speedfam have a data sheet that travel 
          with them containing cut rate and serial number.  Typical 
          cut rates used here at Digital are 14 to 18, with 16-17 
          being most common so far.
        - The lower the number (ie; 14) the lower the cut rate 
          (removal rate).
31.16New Storage Box For Alignment FixturesASDG::POIRIERSat Mar 01 1997 19:077
    
    We have aquired a new, (larger) pelican box to store the Speedfam
    alignment fixtures in. The new box is grey in color and it is located
    on top of the Speedfam file cabinet next to CMP.B2.
    
    
                                                           C Shift. 
31.17wafer detect testedFABSIX::J_SWENSONSat Mar 15 1997 16:2418
    wafer detect tested on B4
    -------------------------
    
    As requested by engineering a test was performed to check the
    wafer detect system.
    
    The test was performed on B4 running the IOMQC_20 recipe.
    
    Five wafers were loaded and ran, during segment 5 the wafer
    detect sensor was tripped.
    
    Results
    
    -Tool alarmed and table stopped immediately
    -All carriers raised up
    
    
                                       C-shift
31.18MEGACROPOLIS BATMANFABSIX::M_RIDLONEliminate the obliviousTue Mar 25 1997 06:426
    MEGA SLURRY DELIVERY ALARM IS NOW TIED INTO B2 AND B4. IT HAS BEEN
    TESTED AND WORKS. OTHER TOOLS WILL BE WIRED UP WHEN AVAILABLE. 
    
    
    
    						B-SHIFT
31.19ionizers installed todayFABSIX::J_SWENSONFri Mar 28 1997 19:253
    Ionizers have been installed in B1,B2 and B3 today.
    
                                 C-shift
31.20INFRANORS NEED TO BE PICKED UP.FABSIX::B_WESTPHALSat Apr 05 1997 20:258
General note - Infranors were to be shipped to Packaging Plus down the street.
	 John S. checked for the parts at 10:30 and 11:30-noon (closing time)
	 but they never showed up.  Be aware that none are in stock and some 
	 one needs to pick them up on Monday.